# Calderon-preconditioned boundary integral equations of the Burton–Miller type for transmission problems

### v1

@@ 5 The present formulation: Multi-material cases
Similarly to (20), it is easy to see that (65) is also highly ill-conditioned. If we naively extend the integral formalism (45) to the current problem, we multiply the complex constant $\beta=-\alpha$ to the last three equations in (52) and “appropriately” arrange the equations, which gives the following:
@@
${\cal A}_{\beta;\mathrm{orig}}^{2}=\begin{pmatrix}\frac{\beta(\beta-\alpha\varepsilon_{2})}{4}{\cal I}&0&0&0\\ -\frac{\alpha\beta\varepsilon_{2}}{4}{\cal I}&\frac{\beta^{2}}{4}{\cal I}&0&0\\ \frac{\beta+2\alpha}{4}{\cal I}+\frac{\alpha(\beta+\alpha)}{2}{\cal N}_{1}-\alpha(\beta{\cal N}_{1}{\cal D}_{2}+\alpha{\cal D}_{1}^{*}{\cal N}_{1})&-\alpha\beta{\cal N}_{1}{\cal D}_{2}&\frac{\alpha(\alpha-\beta\varepsilon_{2})}{4}{\cal I}&-\frac{\alpha\beta\varepsilon_{2}}{4}{\cal I}\\ 0&\frac{\beta^{2}}{4}{\cal I}&0&0\end{pmatrix}+{\cal K},$ (83)
with a compact operator ${\cal K}$. Unfortunately, the right-bottom block of (83) ends up with a zero operator regardless of the choice of $\beta$. Due to the zero block in the diagonal, the eigenvalue of ${\cal A}^{2}_{\beta;\mathrm{orig}}$ may cluster at zero in the complex plane, yielding a possibly large condition number after the discretisation.
@@
$\begin{pmatrix}\beta\left(\frac{1}{2}{\cal I}-{\cal D}_{1}\right)&-\beta{\cal D}_{2}&\beta\varepsilon_{2}{\cal S}_{2}&\beta\varepsilon_{2}{\cal S}_{2}\\ -\beta{\cal D}_{2}&\beta\left(\frac{1}{2}{\cal I}-{\cal D}_{1}\right)&\beta\varepsilon_{2}{\cal S}_{2}&\beta\varepsilon_{2}{\cal S}_{2}\\ \frac{1}{2}{\cal I}+{\cal D}_{1}+\alpha{\cal N}_{1}&0&\alpha\left(\frac{1}{2}{\cal I}-{\cal D}_{1}^{*}\right)-{\cal S}_{1}&0\\ 0&\frac{1}{2}{\cal I}+{\cal D}_{3}+\gamma{\cal N}_{3}&0&\gamma\varepsilon_{3}\left(\frac{1}{2}{\cal I}-{\cal D}_{3}^{*}\right)-\varepsilon_{3}{\cal S}_{3}&\end{pmatrix}\begin{pmatrix}u_{a}\\ u_{b}\\ w_{a}\\ w_{b}\\ \end{pmatrix}=\begin{pmatrix}0\\ 0\\ u^{\mathrm{in}}_{a}+\alpha q^{\mathrm{in}}_{a}\\ 0\end{pmatrix},$ (96)
@@
${\cal A}_{-\alpha,\alpha/\varepsilon_{3};\mathrm{mod}}^{2}=\begin{pmatrix}\frac{\alpha^{2}(1+\varepsilon_{2})}{4}{\cal I}&\frac{\alpha^{2}\varepsilon_{2}}{4\varepsilon_{3}}{\cal I}&0&0\\ \frac{\alpha^{2}\varepsilon_{2}}{4}{\cal I}&\frac{\alpha^{2}\left(1+\frac{\varepsilon_{2}}{\varepsilon_{3}}\right)}{4}{\cal I}&0&0\\ \frac{\alpha}{4}{\cal I}&\alpha^{2}{\cal N}_{1}{\cal D}_{2}&\frac{\alpha^{2}(1+\varepsilon_{2})}{4}{\cal I}&\frac{\alpha^{2}\varepsilon_{2}}{4}{\cal I}\\ \frac{\alpha^{2}}{\varepsilon_{3}}{\cal N}_{3}{\cal D}_{2}&\frac{\alpha}{4}{\cal I}&\frac{\alpha^{2}\varepsilon_{2}}{4\varepsilon_{3}}{\cal I}&\frac{\alpha^{2}\left(1+\frac{\varepsilon_{2}}{\varepsilon_{3}}\right)}{4}{\cal I}\end{pmatrix}+{\cal K}.$ (101)
Note that, by the parameter setting, we could vanish the hypersingular operators in $(3,1)$ and $(4,2)$ blocks of (101). Although it is impossible to eliminate all the hypersingular operators, ${\cal A}^{2}_{-\alpha,\alpha/\varepsilon_{3};\mathrm{mod}}$ might be well-conditioned. Since the non-compact part of ${\cal A}^{2}_{-\alpha,\alpha/\varepsilon_{3};\mathrm{mod}}$ has the lower triangular block structure, the hypersingular operators in the lower-left blocks do not affect the eigenvalue accumulation points of the whole operator. It is easy to see that the eigenvalue clusters are found only at $\frac{\alpha^{2}}{4}$ and $\frac{\alpha^{2}(\varepsilon_{2}+\varepsilon_{3}+\varepsilon_{2}\varepsilon_{3})}{4\varepsilon_{3}}\neq 0$. We may thus conclude that the proposed integral formalism (96) with $\beta=-\alpha$ and $\gamma=\alpha/\varepsilon_{3}$ is well-conditioned. Note that there is another possibility as the integral equations (96); The equations in the second and fourth rows can be interchanged. In that case, one needs to use the ordinary and the Burton–Miller formulations for the new second and fourth rows, respectively, to have a well-conditioned system. The parameters $\beta$ and $\gamma$ should also be adjusted so that as many as the hypersingular operators are eliminated from the square of the underlying operator.
@@ 6 Numerical demonstrations
In this section, we demonstrate the validity and efficiency of the proposed integral formalism through several numerical experiments. The boundary integral equations are discretised by the collocation method with constant elements. The boundary integrals are evaluated by the Gauss–Legendre quadrature with 10 Gaussian nodes. In the evaluation, the weak singularity for (9) is handled with a standard “subtraction and addition” technique, and the integral with the hypersingular kernel (12) is also regularised. As the linear algebraic solver, the GMRES without restart is utilised, and its tolerance is set as $10^{-8}$.
@@ 6.1 Single material cases
We then verify the efficiency dependence of our formulation on excitation angular frequency $\omega$ and the contrast ratio in material constants $\varepsilon_{2}~{}(=\varepsilon_{2}/\varepsilon_{1})$ of the inclusion and the exterior medium. To this end, we again use the circular example (Figure 2) and measure the number of GMRES iterations for various $\omega$ and $\varepsilon_{2}$. The left and right of Figure 5 show the number of iterations vs $\omega$ and $\varepsilon_{2}$, respectively. In the computation for the left figure, $\varepsilon_{2}$ is fixed as $2.0$. For right, $\omega$ is fixed as $5.0$. For both computations, the number of boundary elements $N$ is fixed as $N=600$. The same observations as the previous example are also seen in these examples; The proposed formulations as well as the PMCHWT and SBIEs outperform the conventional Burton–Miller method, and the proposed one with $\beta=-\alpha$ is almost always faster than the one with $\beta=1$, and the performance of the proposed methods slightly worse than the PMCHWT and SBIEs. The performance difference is, however, not significant. It is also worth noting that, in Figure 5, while there are several peaks in the plot for the conventional Burton–Miller method, there are none for the rest of the methods. The peaks correspond to the resonance angular frequencies; At these frequencies, the transmitted wave is trapped in the inclusion $\Omega_{2}$. It is well-known that the linear system obtained from the boundary integral equation may considerably be ill-conditioned for such a case. Nonetheless, the proposed formulations (as well as the PMCHWT and SBIEs) robustly work even in such a case. With these observations, especially given the high accuracy (see Figure 3), the proposed method would also be attractive when solving transmission problems consisting of a single scatterer.
@@ 6.2 Multi-material cases
We then check the performance dependence of the proposed methods on the incident angular frequency $\omega$ and the material constants $\varepsilon_{2}$ and $\varepsilon_{3}$. For the former, we fix the material constants as before: $\varepsilon_{i}=i$ for $i=1,\cdots,3$. For the latter, we fix $\omega=5.0$ and sweep the inclusion material constants as $\varepsilon_{2}=1/\varepsilon_{3}$ for $0<\varepsilon_{3}\leq 10$. For both computations, the inner and outer circular surfaces are discretised into 600 and 1200 constant boundary elements, respectively. Figure 11 show the number of GMRES iterations for conventional and proposed BEMs of the Burton–Miller type.
@@
The eigenvalues of the square of the proposed coefficint matrices cluster around the origin, while those of the conventional formulation diverge. These spectral properties account for the fast GMRES convergence of the proposed formulations. The right of Figure 12 shows the close-up near the origin of the left one. One finds that the eigenvalues for the proposed formulations cluster at a few points. Note here that not all the eigenvalues for ($\beta=-\alpha$) and (mod.) are plotted in the right figure. The range is truncated so that one easily understands the behaviour of the eigenvalues in the vicinity of their clustering points. The eigenvalues with the largest magnitude for ($\beta=-\alpha$) and (mod.) are found at $2.1706-2.4538\mathrm{i}$ and $-5.9178+3.9583\mathrm{i}$, respectively. In the right figure (the magnified one), the theoretically estimated accumulation points for the integral operators are also plotted. The actual clustering points are a bit different from the estimated position. This inconsistency is naturally expected from the discussions for the single-material case; see Figures 7 and 8. The hypersingular operators in (83) and (101) are responsible for the inconsistency. As expected, the original version of the Calderon-preconditioned equation with $\beta=-\alpha$ involves the coefficient matrix whose square has eigenvalues with a tiny magnitude. On the other hand, the coefficient matrix of the modified version does not. These spectral properties explain the performance difference of the proposed solvers, concluding that the modified version is superior to the original one.
@@ 7 Conclusions
In this paper, we proposed novel boundary integral formulations based on the Burton–Miller formulation tailored for the transmission problem. We demonstrated that (1) appropriately ordering the integral operators and (2) multiplying proper constants for some rows and/or replacing some of the operators by the Burton–Miller ones can provide a well-conditioned integral formalism. Specifically, the square of the whole integral operator can be made equal to the one with a few eigenvalues except for a compact perturbation, providing a fast convergence for GMRES applied to the algebraic equations obtained by discretising the integral equation. Even when not all the hypersingular operators are eliminated from the square, fast convergence is achieved. Our formulation is easy to implement and can systematically be applied to various problems such as scattering by an object consisting of several materials with “material junction points”.
@@
Author contributions
Y. M. conceptualised the study, specifically developed the idea of rearranging the integral operators and found the best choice of the parameter $\beta$ for single-material case. A. Y. implemented test codes, corrected data, and drafted the manuscript. H. I. also conceptualised the study, specifically modifying Y. M.’s original formulation with the Calderon formulae and developed the multi-material formulations, and contributed to drafting and editing the manuscript while supervising the study.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Numbers 23H03413 and 23K19972.

### v2

@@ 6 The present formulation: Multi-material cases
Similarly to (14), it is easy to see that (25) is highly ill-conditioned. If we naively extend the integral formalism (21) to the current problem, we multiply the complex constant $\beta=-\alpha$ to the last three equations in (24) and “appropriately” arrange the equations, which gives the following:
@@
${\cal A}_{\beta;\mathrm{orig}}^{2}=\begin{pmatrix}\frac{\beta(\beta-\alpha\varepsilon_{2})}{4}{\cal I}&0&0&0\\ {\color[rgb]{0,0,0}0}&\frac{\beta^{2}}{4}{\cal I}&0&0\\ \frac{\beta+2\alpha}{4}{\cal I}+\frac{\alpha(\beta+\alpha)}{2}{\cal N}_{1}-\alpha(\beta{\cal N}_{1}{\cal D}_{2}+\alpha{\cal D}_{1}^{*}{\cal N}_{1})&{\color[rgb]{0,0,0}0}&\frac{\alpha(\alpha-\beta\varepsilon_{2})}{4}{\cal I}&{\color[rgb]{0,0,0}0}&\\ 0&\frac{\beta^{2}}{4}{\cal I}&0&0\end{pmatrix}+{\cal K},$ (27)
with a compact operator ${\cal K}$. Unfortunately, the right-bottom block of (27) ends up with a zero operator regardless of the choice of $\beta$. Due to the zero block in the diagonal and the block diagonal structure, the eigenvalue of ${\cal A}^{2}_{\beta;\mathrm{orig}}$ may cluster at zero in the complex plane, yielding a possibly large condition number after the discretisation via collocation.
@@
$\begin{pmatrix}\beta\left(\frac{1}{2}{\cal I}-{\cal D}_{2}\right)&-\beta{\cal D}_{2}&\beta\varepsilon_{2}{\cal S}_{2}&\beta\varepsilon_{2}{\cal S}_{2}\\ -\beta{\cal D}_{2}&\beta\left(\frac{1}{2}{\cal I}-{\cal D}_{2}\right)&\beta\varepsilon_{2}{\cal S}_{2}&\beta\varepsilon_{2}{\cal S}_{2}\\ \frac{1}{2}{\cal I}+{\cal D}_{1}+\alpha{\cal N}_{1}&0&\alpha\left(\frac{1}{2}{\cal I}-{\cal D}_{1}^{*}\right)-{\cal S}_{1}&0\\ 0&\frac{1}{2}{\cal I}+{\cal D}_{3}+\gamma{\cal N}_{3}&0&\gamma\varepsilon_{3}\left(\frac{1}{2}{\cal I}-{\cal D}_{3}^{*}\right)-\varepsilon_{3}{\cal S}_{3}&\end{pmatrix}\begin{pmatrix}u_{a}\\ u_{b}\\ w_{a}\\ w_{b}\\ \end{pmatrix}=\begin{pmatrix}0\\ 0\\ u^{\mathrm{in}}_{a}+\alpha q^{\mathrm{in}}_{a}\\ 0\end{pmatrix},$ (28)
@@
${\cal A}_{-\alpha,\alpha/\varepsilon_{3};\mathrm{mod}}^{2}=\begin{pmatrix}\frac{\alpha^{2}(1+\varepsilon_{2})}{4}{\cal I}&{\color[rgb]{0,0,0}0}&0&0\\ {\color[rgb]{0,0,0}0}&\frac{\alpha^{2}\left(1+\frac{\varepsilon_{2}}{\varepsilon_{3}}\right)}{4}{\cal I}&0&0\\ \frac{\alpha}{4}{\cal I}&{\color[rgb]{0,0,0}0}&\frac{\alpha^{2}(1+\varepsilon_{2})}{4}{\cal I}&{\color[rgb]{0,0,0}0}\\ {\color[rgb]{0,0,0}0}&\frac{\alpha}{4}{\cal I}&{\color[rgb]{0,0,0}0}&\frac{\alpha^{2}\left(1+\frac{\varepsilon_{2}}{\varepsilon_{3}}\right)}{4}{\cal I}\end{pmatrix}+{\cal K}.$ (29)
It is easy to see that the eigenvalue clusters are found only at $\frac{\alpha^{2}(1+\varepsilon_{2})}{4}$ and $\frac{\alpha^{2}(1+\varepsilon_{2}/\varepsilon_{3})}{4}\neq 0$. We may thus conclude that the proposed integral formalism (28) with $\beta=-\alpha$ and $\gamma=\alpha/\varepsilon_{3}$ is well-conditioned. Note that there is another possibility as the integral equations (28); The equations in the second and fourth rows can be interchanged, for example. In that case, one needs to use the ordinary and the Burton–Miller formulations for the new second and fourth rows, respectively, to have a well-conditioned system. The parameters $\beta$ and $\gamma$ should also be adjusted so that all the hypersingular operators are eliminated from the square of the underlying operator.
@@ 7 Numerical demonstrations
In this section, we demonstrate the validity and efficiency of the proposed integral formalism through several numerical experiments. The boundary integral equations are discretised by the collocation method with constant elements. The boundary integrals are evaluated by the Gauss–Legendre quadrature with 10 Gaussian nodes. In the evaluation, the weak singularity for (9) is handled with a standard “subtraction and addition” technique, and the integral with the hypersingular kernel (12) is also regularised [29]. As the linear algebraic solver, the GMRES without restart is utilised, and its tolerance is set as $10^{-8}$.
@@ 7.1 Single material cases
We then verify the efficiency dependence of our formulation on excitation angular frequency $\omega$ and the contrast ratio in material constants $\varepsilon_{2}~{}(=\varepsilon_{2}/\varepsilon_{1})$ of the inclusion and the exterior medium. To this end, we again use the circular example (Figure 2) and measure the number of GMRES iterations for various $\omega$ and $\varepsilon_{2}$. The left and right of Figure 5 shows the number of iterations vs $\omega$ and $\varepsilon_{2}$, respectively. In the computation for the left figure, $\varepsilon_{2}$ is fixed as $2.0$. For right, $\omega$ is fixed as $5.0$. For both computations, the number of boundary elements $N$ is fixed as $N=600$. The same observations as the previous example are also seen in these examples; The proposed formulations as well as the PMCHWT and SBIEs outperform the conventional Burton–Miller method, and the proposed one with $\beta=-\alpha$ is almost always faster than the one with $\beta=1$, and the performance of the proposed methods slightly worse than the PMCHWT and SBIEs. The performance difference is, however, not significant. It is also worth noting that, in Figure 5, while there are several peaks in the plot for the conventional Burton–Miller method, there are none for the rest of the methods. The peaks correspond to the resonance angular frequencies; At these frequencies, the transmitted wave is trapped in the inclusion $\Omega_{2}$. It is well-known that the linear system obtained from the boundary integral equation may considerably be ill-conditioned for such a case. Nonetheless, the proposed formulations (as well as the PMCHWT and SBIEs) robustly work even in such a case. With these observations, especially given the high accuracy (see Figure 3), the proposed method would also be attractive when solving transmission problems consisting of a single scatterer.
@@ 7.2 Multi-material cases
We then check the performance dependence of the proposed methods on the incident angular frequency $\omega$ and the material constants $\varepsilon_{2}$ and $\varepsilon_{3}$. For the former, we fix the material constants as before: $\varepsilon_{i}=i$ for $i=1,\cdots,3$. For the latter, we fix $\omega=5.0$ and sweep the inclusion material constants as $\varepsilon_{2}=1/\varepsilon_{3}$ for $0<\varepsilon_{3}\leq 10$. For both computations, the inner and outer circular surfaces are discretised into 600 and 1200 constant boundary elements, respectively. Figure 11 shows the number of GMRES iterations for conventional and proposed BEMs of the Burton–Miller type.
@@
The eigenvalues of the square of the proposed coefficint matrices cluster around the origin, while those of the conventional formulation diverge. These spectral properties account for the fast GMRES convergence of the proposed formulations. The right of Figure 12 shows the close-up near the origin of the left one. One finds that the eigenvalues for the proposed formulations cluster at a few points. Note here that not all the eigenvalues for ($\beta=-\alpha$) and (mod.) are plotted in the right figure. The range is truncated so that one easily understands the behaviour of the eigenvalues in the vicinity of their clustering points. The eigenvalues with the largest magnitude for ($\beta=-\alpha$) and (mod.) are found at $2.1706-2.4538\mathrm{i}$ and $-5.9178+3.9583\mathrm{i}$, respectively. In the right figure (the magnified one), the theoretically estimated accumulation points for the integral operators are also plotted. One confirms that the eigenvalues accumulate around the estimated points. As expected, the original version of the Calderon-preconditioned equation with $\beta=-\alpha$ involves the coefficient matrix whose square has eigenvalues with a tiny magnitude. On the other hand, the coefficient matrix of the modified version does not. These spectral properties explain the performance difference of the proposed solvers, concluding that the modified version is superior to the original one.
@@ 8 Conclusions
In this paper, we proposed novel boundary integral formulations based on the Burton–Miller formulation tailored for the transmission problem. We demonstrated that appropriately ordering the integral operators and multiplying proper constants for some rows and/or replacing some of the operators by the Burton–Miller ones can provide a well-conditioned integral formalism. Specifically, the square of the whole integral operator can be made equal to the one with a few eigenvalue clustrering points, providing a fast convergence for GMRES applied to the algebraic equations obtained by discretising the integral equation via collocation. Our formulation is easy to implement and can systematically be applied to various problems such as scattering by an object consisting of several materials with “material junction points”.
@@ Acknowledgments
The authors benefited from discussions with Mr. K. Tomoyasu from Keio University on the multi-material formulations. This work was supported by JSPS KAKENHI Grant Numbers 23H03413, 23K19972 and 24K20783. This work used the computational resources of the supercomputer Camphor3 provided by ACCMS of Kyoto University through the Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures and High Performance Computing Infrastructure in Japan (Project IDs: jh240031).